A 0.115 M solution of a weak acid (HA) has a pH of 3.32. Calculate the acid ionization constant (Ka) for the acid.
Answer #1

[H+] = 10^-3.32 = 0.000478630092 M



HA <-----> H+ + A-
[H+] = [A-]
[HA] = 0.115 - 0.000478630092 = 0.11452137

K = ( 0.0004786)^2 / 0.115 = 1.99180835 × 10-6